OpenStudy (anonymous):

When does a cosine function have 2 solutions?

OpenStudy (anonymous):

when you restrict the domain if you mean solve \[\cos(\alpha)=y\] with \(-1\leq y\leq 1\)

OpenStudy (anonymous):

wait, i mean like 2 solutions for a value of y

OpenStudy (anonymous):

Well, cos(something) = 1 when that something is either zero degrees or 180 degrees.

OpenStudy (anonymous):

can you explain what you mean?

OpenStudy (callisto):

When k = -1, alpha = 180 degrees (or pi radians) ... There is no other angle giving the same value for 1 period. I meant this: